Now, let’s look at a real-world example of economic growth. We want to see the rate at which Peru’s GDP per capita income grew between 1990 and 2018. GDP per capita in 1990 is 2650.214 and GDP per capita in 2018 is 6453.561. Assuming a constant growth rate over this period, what rate has Peru grown at? HINT: use the constant growth rate formula from lecture and discussion: y(t) = y(0) [1+g*]t
Answer #1

ANSWER:

Y(T) = 6,453.561

Y(0) = 2,650.214

T = 2018 - 1990 = 28

6,453.561 = 2,650.214 * (1 + G) ^ 28

6,453.561 / 2,650.214 = (1 + G) ^ 28

2.4351 = (1 + g) ^ 28

(2.4351) ^ 1 / 28 = 1 + g

1.0322 = 1 + g

g = 1.0322 - 1

g = 0.0322 or 3.22%

so the constant growth rate is 3.22%